OpenStudy (anonymous):

can someone please tell me how to use properties [in math] please just describe it [no numbers]

OpenStudy (anonymous):

what kind of properties

OpenStudy (anonymous):

like simplifying equations

OpenStudy (anonymous):

ok give me an equation and ill try to simplify it

OpenStudy (anonymous):

but i need like [how to put this] just words not lsomething like "6 (4+2) =" or whatever

OpenStudy (anonymous):

so i can describe it in detail

OpenStudy (anonymous):

ok lets first talk about the distributive property

OpenStudy (anonymous):

in an equation when it has 4(5x+3) you hav to multiply the two numbers including the x

OpenStudy (anonymous):

sooo........ 4*5x=(20x and 4*3=12 so the answer is 20x+12

OpenStudy (anonymous):

but how to describe that in detail [of like words xD no numbers}

OpenStudy (anonymous):

oh ok ill try

OpenStudy (anonymous):

what kind of property u want me to explain

OpenStudy (anonymous):

Just how to use any property

OpenStudy (anonymous):

ok ill explain the commutative property of addition bc its one of the easiest

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the commutative property of addition is very simple <----- dont write tht down tho :) ok First of all think of any two numbers ok ..................... now add them together once you got your answer add them again but this time switch the numbers this basically is the commutative property which is any number added to another number equals the same number and order doesnt count that was hard with no numbers lol

OpenStudy (anonymous):

so 5+6 equals 11 and 6+5 =11 (lol soory had to include numbers so u would understand)

OpenStudy (anonymous):

lol its k